OpenStudy (anonymous):

find -x when x=77 please explain how I figure this out

OpenStudy (mertsj):

-x means the opposite of x. So if x is 77, then the opposite of x is -77.

OpenStudy (mertsj):

Or you could write this equation: x = 77 Now multiply both sides by negative 1 -1(x) = -1(77) -x = -77

OpenStudy (karatechopper):

-x is -1 ttimes x. meaning -1x=77 and if i divide -1 on both sides then x=-77